OpenStudy (atomicreaper):

Consider the equations y = |x − 1| and y = 3x + 2. The approximate solution of this system of equations is _______________?

OpenStudy (mathmale):

An accurate graph might be all you need to solve this problem. Graph y=3x+2 first. As for the other equation, the absolute value operation makes graphing this just a bit trickier. Keep in mind that y=|x-1| represents two separate equations: y=x-1 and y=-(x-1). The value of x at which one graph becomes the other is x=1.

OpenStudy (mathmale):

Alternatively, graph y=|x|. Then move the entire graph of y=\X\ one unit to the right. This action is called horizontal translation.

OpenStudy (dayakar):

\[y=\left| x-1 \right|---------(1)\] y= 3x+2----------(2) \[\left| x-1 \right| = 3x +2\] [since (1) = (2) can u solve remaining part

OpenStudy (atomicreaper):

i got the points (0,2) and (1,2)

OpenStudy (dayakar):

can u show ur work

OpenStudy (atomicreaper):

No, my computer had to reset and that is where I done it... the site im doing it on kept the answer I put in though.